# A Multi-level Analysis of Factors Associated with Student Performance: A Machine Learning Approach to the SAEB Microdata

### v1

@@ Abstract
Identifying the factors that influence student performance in basic education is a central challenge for formulating effective public policies in Brazil. This study introduces a multi-level machine learning approach to classify the proficiency of 9th-grade and high school students using microdata from the System of Assessment of Basic Education (SAEB). Our model uniquely integrates four data sources: student socioeconomic characteristics, teacher professional profiles, school indicators, and director management profiles. A comparative analysis of four ensemble algorithms confirmed the superiority of a Random Forest model, which achieved 90.2% accuracy and an Area Under the Curve (AUC) of 96.7%. To move beyond prediction, we applied Explainable AI (XAI) using SHAP, which revealed that the school’s average socioeconomic level is the most dominant predictor, demonstrating that systemic factors have a greater impact than individual characteristics in isolation. The primary conclusion is that academic performance is a systemic phenomenon deeply tied to the school’s ecosystem. This study provides a data-driven, interpretable tool to inform policies aimed at promoting educational equity by addressing disparities between schools.
@@ 1 Introduction
    We design and implement a novel multi-level predictive model by systematically integrating four distinct data sources from the national SAEB assessment: student socioeconomic profiles, teacher professional data, consolidated school indicators, and director management profiles.
@@ 2 Related Work
Our Study Develops a multi-level prediction model by integrating student, teacher, school, and director data from the national SAEB assessment to identify systemic factors associated with student performance. Addresses the gap by applying a systemic, multi-level machine learning approach to a recent, large-scale national dataset for Brazilian basic education.
@@ 3 Background: The System of Assessment of Basic Education (SAEB)
The System of Assessment of Basic Education (SAEB) is a comprehensive, large-scale evaluation conducted periodically by the National Institute for Educational Studies and Research Anísio Teixeira (INEP), an agency of the Brazilian Ministry of Education. Its primary objective is to assess the quality of basic education in Brazil and identify factors that may impact student performance [28]. SAEB comprises standardized tests in core subjects, such as Portuguese and Mathematics, and a series of contextual questionnaires administered to students, teachers, school directors, and education secretaries. The resulting microdata provide a rich, multidimensional snapshot of the national educational landscape, making them an invaluable resource for research and policy formulation. For this study, we utilized the microdata from the SAEB administration, focusing on 9th-grade and high school students. The multi-level nature of our analysis was made possible by integrating four distinct, yet interconnected, data sources provided by INEP [34]:
@@
  • DIRECTOR: The dataset contains information from the school director’s questionnaire. It provides insights into the school’s management and leadership, including the director’s professional experience, training in school administration, and community and parental engagement practices.
@@ 4.1 Data Source and Acquisition
The primary data source was the official microdata from the System of Assessment of Basic Education (SAEB), publicly available through the National Institute for Educational Studies and Research Anísio Teixeira (INEP) [34]. To construct a comprehensive multi-level model, this study integrated data from both 9th-grade and high school students, using four distinct datasets: STUDENT, TEACHER, SCHOOL, and DIRECTOR.
@@
Total Unique Directors 73,595
@@
4:  $student\_data,teacher\_data,school\_data,director\_data\leftarrow\textsc{LoadCSVs}(file\_paths)$
5:  $merged\_data\leftarrow\textsc{JoinSources}(student\_data,teacher\_data,school\_data,director\_data)$ $\triangleright$ 2. Data Cleaning and Preprocessing
@@ 4.5 Feature Selection using the Boruta Algorithm
The algorithm was applied to an initial set of variables from the student, teacher, school, and director levels. The results of this analysis were decisive: out of 98 potential features, Boruta confirmed 44 as statistically relevant. The confirmed features were overwhelmingly concentrated in the student and school levels, including variables related to parental education, home resources, and the school’s overall socioeconomic context.
@@ 5.2 Feature Importance Analysis
Director’s Experience School Director’s Years of Experience. More experienced leadership is associated with better school and student outcomes.
@@ 5.4.2 Implications for School Leaders
For school directors and pedagogical coordinators, the SHAP explanations enable the development of targeted, student-level interventions.
@@ Appendix A Variable Glossary
Director Variables (TS_DIRETOR)
TX_Q014 Nível de escolaridade Director’s Level of Education What is the HIGHEST level of education the director has achieved.
TX_Q016 Tempo de Experiência como Diretor Director’s Years of Experience The director’s total years of experience in a leadership position.

### v2

@@ Abstract
Identifying the factors that influence student performance in basic education is a central challenge for formulating effective public policies in Brazil. This study introduces a multi-level machine learning approach to classify the proficiency of 9th-grade and high school students using microdata from the System of Assessment of Basic Education (SAEB). Our model uniquely integrates four data sources: student socioeconomic characteristics, teacher professional profiles, school indicators, and principal management profiles. A comparative analysis of four ensemble algorithms confirmed the superiority of a Random Forest model, which achieved 90.2% accuracy and an Area Under the Curve (AUC) of 96.7%. To move beyond prediction, we applied Explainable AI (XAI) using SHAP, which revealed that the school’s average socioeconomic level is the most dominant predictor, demonstrating that systemic factors have a greater impact than individual characteristics in isolation. The primary conclusion is that academic performance is a systemic phenomenon deeply tied to the school’s ecosystem. This study provides a data-driven, interpretable tool to inform policies aimed at promoting educational equity by addressing disparities between schools.
@@ 1 Introduction
    We design and implement a novel multi-level predictive model by systematically integrating four distinct data sources from the national SAEB assessment: student socioeconomic profiles, teacher professional data, consolidated school indicators, and principal management profiles.
@@ 2 Related Work
Our Study Develops a multi-level prediction model by integrating student, teacher, school, and principal data from the national SAEB assessment to identify systemic factors associated with student performance. Addresses the gap by applying a systemic, multi-level machine learning approach to a recent, large-scale national dataset for Brazilian basic education.
@@ 3 Background: The System of Assessment of Basic Education (SAEB)
The System of Assessment of Basic Education (SAEB) is a comprehensive, large-scale evaluation conducted periodically by the National Institute for Educational Studies and Research Anísio Teixeira (INEP), an agency of the Brazilian Ministry of Education. Its primary objective is to assess the quality of basic education in Brazil and identify factors that may impact student performance [28]. SAEB comprises standardized tests in core subjects, such as Portuguese and Mathematics, and a series of contextual questionnaires administered to students, teachers, school principals, and education secretaries. The resulting microdata provide a rich, multidimensional snapshot of the national educational landscape, making them an invaluable resource for research and policy formulation. For this study, we utilized the microdata from the SAEB administration, focusing on 9th-grade and high school students. The multi-level nature of our analysis was made possible by integrating four distinct, yet interconnected, data sources provided by INEP [34]:
@@
  • PRINCIPAL: The dataset contains information from the school principal’s questionnaire. It provides insights into the school’s management and leadership, including the principal’s professional experience, training in school administration, and community and parental engagement practices.
@@ 4.1 Data Source and Acquisition
The primary data source was the official microdata from the System of Assessment of Basic Education (SAEB), publicly available through the National Institute for Educational Studies and Research Anísio Teixeira (INEP) [34]. To construct a comprehensive multi-level model, this study integrated data from both 9th-grade and high school students, using four distinct datasets: STUDENT, TEACHER, SCHOOL, and PRINCIPAL.
@@
Total Unique principals 73,595
@@
4:  $student\_data,teacher\_data,school\_data,principal\_data\leftarrow\textsc{LoadCSVs}(file\_paths)$
5:  $merged\_data\leftarrow\textsc{JoinSources}(student\_data,teacher\_data,school\_data,principal\_data)$ $\triangleright$ 2. Data Cleaning and Preprocessing
@@ 4.5 Feature Selection using the Boruta Algorithm
The algorithm was applied to an initial set of variables from the student, teacher, school, and principal levels. The results of this analysis were decisive: out of 98 potential features, Boruta confirmed 44 as statistically relevant. The confirmed features were overwhelmingly concentrated in the student and school levels, including variables related to parental education, home resources, and the school’s overall socioeconomic context.
@@ 5.2 Feature Importance Analysis
Principal’s Experience School Principal’s Years of Experience. More experienced leadership is associated with better school and student outcomes.
@@ 5.4.2 Implications for School Leaders
For school principals and pedagogical coordinators, the SHAP explanations enable the development of targeted, student-level interventions.
@@ Appendix A Variable Glossary
Principal Variables (TS_DIRETOR)
TX_Q014 Nível de escolaridade Principal’s Level of Education What is the HIGHEST level of education the principal has achieved.
TX_Q016 Tempo de Experiência como Diretor Principal’s Years of Experience The principal’s total years of experience in a leadership position.